Circular No. 7398

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000bi IN ANONYMOUS GALAXY M. M. M. Santangelo, Lucca, Italy, reports his discovery with the

Monte Agliale 0.51-m reflector, in the course of the Monte Agliale Supernovae Search, of an apparent supernova in a member galaxy of the Abell 576 cluster. SN 2000bi was found at R ∼ 18.4 on three unfiltered frames taken on Mar. 22.90 UT, and confirmed on ten frames taken on Apr. 7.92 (R ∼ 19.2) and 8.85 (R ∼ 19.2); the limiting magnitude of all these frames is R ∼ 20.0. The new object is located at α = 7^{h}21^{m}33^{s}.46, δ = +55^{o}45′30′′.9 (equinox 2000.0), which is ∼ 12′′ east and 4′′ north of the host galaxy nucleus. Inspection of first and second Palomar Sky Survey images and of the Palomar Quick-V survey, as well as a frame taken with the 1.52-m reflector of the Loiano Observatory some years ago, reveals nothing at the position of SN 2000bi.

SUPERNOVA 2000bj IN ANONYMOUS GALAXY D. Kirkman, Bell Laboratories (BL); I. Dell’Antonio and R. A. Schom-

mer, National Optical Astronomy Observatory; G. Squires, California Institute of Technology; and J. A. Tyson, BL, of the Deep Lens Survey (see http://dls.bell-labs.com) team, with support from D. Wittman (BL), A. Becker and C. Stubbs (University of Washington), and G. Bernstein (University of Michigan), report the discovery of a possible supernova (V ∼ 24.3) on CCD 600-s V frames taken with the Cerro Tololo 4-m telescope (+ Mosaic II camera) on Apr. 5.039 UT. SN 2000bj is located at α = 10^{h}51^{m}27^{s}.24, δ = −5^{o}08′29′′.6 (equinox 2000.0), which is ∼ 0′′.9 north of the centroid of a faint galaxy. SN 2000bj was also present on subsequent B-, R-, and z-band frames taken on the same night, and in V -band images taken on Apr. 2 (when it was 0.2 mag fainter), but it is not visible on images taken with the same telescope on Mar. 8 (V > 24.7). The lowsurface-brightness host galaxy (V ∼ 24.5), visible in some of the images, is probably a spiral or irregular galaxy. Details and finding charts are posted

at http://www.astro.washington.edu/becker/deeplens/trans4.html.

SUPERNOVA 1999gi IN NGC 3184 Unfiltered CCD magnitudes: Mar. 13.76 UT, 14.9 ± 0.1 (K. Sárneczky,

K. Sziládi, and L. Kiss, Piszkéstető station, Konkoly Observatory); 28.874, 14.9 (K. Hornoch, Lelekovice, Czech Republic); Apr. 2.908, 15.0 (Hornoch); 8.815, 15.3 (Hornoch); 9.816, 15.3 (Hornoch).

2000 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green